You worked 18 hours last week and earned a total of $124 before taxes. Your job as a lifeguard pays $8 per hour, and your job as a cashier pays $6 per hour. How many hours did you work at each job?

Answer by ankor@dixie-net.com(22740)

You worked 18 hours last week and earned a total of $124 before taxes.
Your job as a lifeguard pays $8 per hour, and your job as a cashier pays $6 per hour.
How many hours did you work at each job?
:
Let x = no. of hrs lifeguarding
Let y = no. of hrs cashiering
:
Total hrs equation
x + y = 18
y = (18-x)
;
Total$ equation
8x + 6y = 124
:
Substitute (18-x) for y
8x + 6(18-x) = 124
8x + 108 - 6x = 124
8x - 6x = 124 - 108
2x = 16
x = 8 hrs lifeguarding
and
18-8 = 10 hrs cashiering
:
:
Check: 8(8) + 6(10) = 124